Solve the following system of equations graphically.
y = 2x - 3
x + y = 3
What is the solution set?

Answer by Cromlix(4381)

Hi there,
y = 2x - 3
x + y = 3
Sort out equations:
y - 2x = -3.......(1)
y + x = 3........(2)
Multiply (2) by 2
y - 2x = -3.......(1)
2y + 2x = 6........(2)
Add (1) + (2)
3y = 3
y = 1
Substitute y = 1 into (1)
y - 2x = -3
1 - 2x = -3
-2x = -3 -1
-2x = -4
x = -4/-2
x = 2
Solution set {2,1}
Hope this helps:-)